OpenStudy (anonymous):

What is the solution of -2x + 8 = 20?

OpenStudy (anonymous):

Subtract the 8 from both sides, then multiply by -2 on both sides

OpenStudy (anonymous):

sorry, divide by -2

OpenStudy (bahrom7893):

-2x = 20-8 -2x = 12 x = -6

OpenStudy (anonymous):

Find g f if f(x) = 2x - 1 and g(x) = x + 2. Answer 2x - 3 2x + 1 2x - 1 2x + 3

OpenStudy (anonymous):

it the g o f sorry

OpenStudy (bahrom7893):

g(f) = (2x-1)+2 = 2x+1

OpenStudy (anonymous):

whats the solution of -(x + 1) - 5 > 10?

OpenStudy (anonymous):

what is the solution of x+5 = 15 and the 2 is part of a fraction 2

OpenStudy (bahrom7893):

geeezz man im not doing your entire hw lol

OpenStudy (anonymous):

i have 200 questions to do I just need help with those to please lol